# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ 4.3 Matrix factorization bound
Given two $m\times n$ matrices $A,B$, we are interested in bounding $\min_{S^{\top}S=I}\|SA-B\|_{F}$ using $\|A^{\top}A-B^{\top}B\|_{F}$. The minimum problem is known as the orthogonal Procrustes problem, solved in (Schoenemann, 1964; Schönemann, 1966). Specifically, the minimum is attained at $S=UV^{\top}$, where $U\Sigma V^{\top}=BA^{\top}$ is its singular value decomposition.
If $m\leq n$ and $\mathrm{rank}(A)=m$, then the following lemma from (Tu et al., 2016) establishes that the distance between $A$ and $B$ is of the same order of $\|A^{\top}A-B^{\top}B\|_{F}$.
Lemma 3 ((Tu et al., 2016, Lemma 5.4)).
@@
=$\overset{(i)}{=}\|A\|_{F}^{2}+\|B\|_{F}^{2}-2\mathrm{tr}(\Sigma)$
@@
=$\overset{(ii)}{=}\;$ $\sum\nolimits_{i=1}^{d}\sigma_{i}(A^{\top}A+B^{\top}B)-2\sum\nolimits_{i=1}^{d^{\prime}}\sigma_{i}(BA^{\top}),$
@@
4.4 Linear regression bound
A standard assumption in analyzing linear regression $y=A^{\ast}x+e$ is that $\mathrm{Cov}(x)$ has full rank. However, as discussed in Section 3.1, we need to handle rank deficient $\mathrm{Cov}(x)$ in the first $\ell$ steps of system identification.
The following lemma is the main result in Section 4.4.
Lemma 5 (Noisy rank deficient linear regression).
Define random vector $y^{\ast}:=A^{\ast}x^{\ast}+e$, where $x^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$ and $e\sim\mathcal{N}(0,\Sigma_{e})$ are $d_{1}$ and $d_{2}$ dimensional Gaussian random vectors, respectively. Define $x:=x^{\ast}+\delta_{x}$ and $y:=y^{\ast}+\delta_{y}$ where the noise vectors $\delta_{x}$ and $\delta_{y}$ can be correlated with $x^{\ast}$ and $y^{\ast}$, and their $\ell_{2}$-norms are sub-Gaussian with $\mathbb{E}[\|\delta^{x}\|]\leq\epsilon_{x}$, $\|\|\delta^{x}\|\|_{\psi_{2}}\leq\epsilon_{x}$, and $\mathbb{E}[\|\delta^{y}\|]\leq\epsilon_{y}$, $\|\|\delta^{y}\|\|_{\psi_{2}}\leq\epsilon_{y}$. Assume that $\|A^{\ast}\|_{2}$, $\|\Sigma_{\ast}\|_{2}$ and $\|\Sigma_{e}\|_{2}$ are $\mathcal{O}(1)$. Let $\sigma_{\min}^{+}(\Sigma_{\ast}^{1/2})\geq\beta>0$. Suppose we get $n$ observations $x^{(i)}$ and $y^{(i)}$ of $x$ and $y$, where $((x^{\ast})^{(i)})_{i=1}^{n}$ are independent and $(\delta_{x}^{(i)},\delta_{y}^{(i)})_{i=1}^{n}$ can be correlated. Assume that $\sigma_{\min}^{+}(\sum_{i=1}^{n}x^{(i)}(x^{(i)})^{\top})\geq\theta^{2}n$, for some $\theta>0$ that satisfies $\theta=\Omega(\epsilon_{x})$, $\theta=\Omega(\epsilon_{y})$ for absolute constants and $\theta$ has at least $n^{-1/4}$ dependence on $n$. Consider the minimum Frobenius norm solution
@@ 4.4 Linear regression bound
Then, there exists an absolute constant $c>0$, such that if $n\geq c(d_{1}+d_{2}+\log(1/p))$, with probability at least $1-p$,
$\|(\hat{A}-A^{\ast})\Sigma_{\ast}^{1/2}\|_{2}=\mathcal{O}((\beta^{-1}\epsilon_{x}+\epsilon_{y})\log^{1/2}(n/p)+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}).$
@@
$\hat{A}DG^{\top}GD^{\top}+\hat{A}(DG^{\top}\Delta_{x}+\Delta_{x}^{m}{\top}GD^{\top}+\Delta_{x}^{\top}\Delta_{x})$
=\;= $A^{\ast}DG^{\top}GD^{\top}+A^{\ast}DG^{\top}\Delta_{x}+(E^{\top}+\Delta_{y}^{\top})(GD^{\top}+\Delta_{x}).$
@@
=\;= $A^{\ast}DG^{\top}\Delta_{x}-\hat{A}(DG^{\top}\Delta_{x}+\Delta_{x}^{\top}GD^{\top}+\Delta_{x}^{\top}\Delta_{x})+(E^{\top}+\Delta_{y}^{\top})(GD^{\top}+\Delta_{x}).$
@@
$\|(\hat{A}-A^{\ast})D\|_{2}$
$=$ $\big{\|}\big{(}A^{\ast}DG^{\top}\Delta_{x}-\hat{A}(DG^{\top}\Delta_{x}+\Delta_{x}^{\top}GD^{\top}+\Delta_{x}^{\top}\Delta_{x})+(E^{\top}+\Delta_{y}^{\top})(GD^{\top}+\Delta_{x})\big{)}(D^{\top})^{\dagger}(G^{\top}G)^{-1}\big{\|}_{2}.$
We claim that as long as $n\geq 16(d_{1}+d_{2}+\log(1/p))$, with probability at least $1-4p$, $\|\hat{A}\|_{2}=\mathcal{O}(\|A^{\ast}\|_{2})$ (Claim 1). Then, since $\|D^{\dagger}\|_{2}=(\sigma_{\min}^{+}(\Sigma_{\ast}))^{-1}\leq\beta^{-1}$,
$\|(\hat{A}-A^{\ast})D\|_{2}$
=\;= $\mathcal{O}(\beta^{-1}\|A^{\ast}\|_{2})\big{(}\|D\|_{2}\|G\|_{2}\|\Delta_{x}\|_{2}+\|\Delta_{x}\|_{2}^{2})\|(G^{\top}G)^{-1}\|_{2}\big{)}$
@@
By (Wainwright, 2019, Theorem 6.1), the Gaussian ensemble $G$ satisfies that with probability at least $1-p$,
@@
[…]
Since $n\geq 8d_{1}+16\log(1/p)$, we have $\|G\|_{2}=\mathcal{O}(n^{1/2})$ and $\sigma_{\min}(G)=\Omega(n^{1/2})$. It follows that $\|(G^{\top}G)^{-1}\|_{2}=\mathcal{O}(n^{-1})$ and $\|G^{\dagger}\|_{2}=\mathcal{O}(n^{-1/2})$. Similarly, $\|E\|_{2}=\mathcal{O}((\|\Sigma_{e}\|_{2}n)^{1/2})$. Note that $\Delta_{x}^{\top}\Delta_{x}=\sum_{i=1}^{n}\delta_{x}^{(i)}(\delta_{x}^{(i)})^{\top}$. Since $\|\delta_{x}^{(i)}\|$ is sub-Gaussian with $\mathbb{E}[\|\delta_{x}^{(i)}\|]\leq\epsilon_{x}$ and $\|\|\delta_{x}^{(i)}\|\|_{\psi_{2}}\leq\epsilon_{x}$, with probability at least $1-p$, $\|\delta_{x}^{(i)}\|=\mathcal{O}(\epsilon_{x}\log^{1/2}(n/p))$. Hence,
$\|\Delta_{x}\|_{2}\leq\|\Delta_{x}\|_{F}=\big{(}\sum\nolimits_{i=1}^{n}\|\delta_{x}^{(i)}\|^{2}\big{)}^{1/2}=\mathcal{O}(\epsilon_{x}n^{1/2}\log^{1/2}(n/p)).$
Similarly, $\|\Delta_{y}\|_{2}=\mathcal{O}(\epsilon_{y}n^{1/2}\log^{1/2}(n/p))$. Hence,
$\|(\hat{A}-A^{\ast})D\|_{2}=\;$ $\mathcal{O}(\beta^{-1}\|A^{\ast}\|_{2})(\|D\|_{2}\epsilon_{x}\log^{1/2}(n/p)+\epsilon_{x}^{2}\log(n/p))$
$\quad+\mathcal{O}(1)(\|G^{\dagger}E\|_{2}+\epsilon_{y}\log^{1/2}(n/p)$
$\qquad+\beta^{-1}(\|\Sigma_{e}^{1/2}\|_{2}+\epsilon_{y}\log^{1/2}(n/p))\epsilon_{x}\log^{1/2}(n/p))$
=\;= $\mathcal{O}(\beta^{-1}\|A^{\ast}\|_{2}\epsilon_{x}\log^{1/2}(n/p)+\epsilon_{y}\log^{1/2}(n/p)+\|G^{\dagger}E\|_{2}),$
where we consider $\epsilon_{x}$ and $\epsilon_{y}$ as quantities much smaller than one such that terms like $\epsilon_{x}^{2},\epsilon_{x}\epsilon_{y}$ are absorbed into $\epsilon_{x},\epsilon_{y}$. It remains to control $\|G^{\dagger}E\|_{2}$. (Mhammedi et al., 2020, Section B.2.11) proves via a covering number argument that with probability at least $1-p$,
@@
=\;= $\mathcal{O}(\beta^{-1}\|A^{\ast}\|_{2}\epsilon_{x}\log^{1/2}(n/p)+\epsilon_{y}\log^{1/2}(n/p)$
$\quad+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}),$
@@
Claim 1.
Under the conditions in Lemma 5, as long as $n\geq 16(d_{1}+d_{2}+\log(1/p))$, with probability at least $1-4p$, $\|\hat{A}\|_{2}=\mathcal{O}(\|A^{\ast}\|_{2})$.
Proof.
A minimum norm solution is given by the following closed-form solution of $\hat{A}$ using pseudoinverse (Moore-Penrose inverse):
$\hat{A}=\;$ $(A^{\ast}(X^{\ast})^{\top}+E^{\top}+\Delta_{y}^{\top})X(X^{\top}X)^{\dagger}$
=\;= $A^{\ast}(X^{\dagger}X^{\ast})^{\top}+(X^{\dagger}E)^{\top}+(X^{\dagger}\Delta_{y})^{\top}$
=\;= $A^{\ast}(X^{\dagger}X)^{\top}-A^{\ast}(X^{\dagger}\Delta_{x})^{\top}+(X^{\dagger}E)^{\top}+(X^{\dagger}\Delta_{y})^{\top}.$
@@
We have shown in the proof of Lemma 5 that with probability at least $1-2p$,
$\|\Delta_{x}\|_{2}=\mathcal{O}(\epsilon_{x}n^{1/2}\log^{1/2}(n/p)),\quad\|\Delta_{y}\|_{2}=\mathcal{O}(\epsilon_{y}n^{1/2}\log^{1/2}(n/p)).$
Similar to the proof of Lemma 5, by (Mhammedi et al., 2020, Section B.2.11), with probability at least $1-p$,
@@
$\|\hat{A}\|_{2}\leq\;$ $\|A^{\ast}\|_{2}+\mathcal{O}(\theta^{-1}n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2})$
$\quad+\mathcal{O}((\|A^{\ast}\|_{2}\epsilon_{x}+\epsilon_{y})n^{1/2}\log^{1/2}(n/p)\theta^{-1}n^{-1/2})$
=\;= $\mathcal{O}(\|A^{\ast}\|_{2}(1+\theta^{-1}\epsilon_{x}\log^{1/2}(n/p))+\theta^{-1}\epsilon_{y}\log^{1/2}(n/p))$
$\quad+\mathcal{O}(\theta^{-1}n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}).$
Hence, as long as $\theta=\Omega(\epsilon_{x})$, $\theta=\Omega(\epsilon_{y})$ for absolute constants and $\theta$ has at least $n^{-1/4}$ dependence on $n$, $\|\hat{A}\|_{2}$ is bounded by $c\|A^{\ast}\|_{2}$ for an absolute constant $c>0$. ∎
In Lemma 5, if $\Sigma_{\ast}$ has full rank and $\sigma_{\min}(\Sigma_{\ast})\geq\beta>0$ and $\sigma_{\min}(\sum_{i=1}^{n}x^{(i)}(x^{(i)})^{\top})=\Omega(\beta^{2}n)$, then
@@
Lemma 6 (Noisy linear regression).
Define random variable $y^{\ast}=A^{\ast}x^{\ast}+e$, where $x^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$ and $e\sim\mathcal{N}(0,\Sigma_{e})$ are $d_{1}$ and $d_{2}$ dimensional random vectors. Assume that $\|A^{\ast}\|_{2}$, $\|\Sigma_{\ast}\|_{2}$ and $\|\Sigma_{e}\|_{2}$ are $\mathcal{O}(1)$, and $\sigma_{\min}(\Sigma_{\ast}^{1/2})\geq\beta>0$. Define $x:=x^{\ast}+\delta_{x}$ and $y:=y^{\ast}+\delta_{y}$ where the noise vectors $\delta_{x}$ and $\delta_{y}$ can be correlated with $x^{\ast}$ and $y^{\ast}$, and their $\ell_{2}$ norms are sub-Gaussian with $\mathbb{E}[\|\delta^{x}\|]\leq\epsilon_{x}$, $\|\|\delta^{x}\|\|_{\psi_{2}}\leq\epsilon_{x}$ and $\mathbb{E}[\|\delta^{y}\|]\leq\epsilon_{y}$, $\|\|\delta^{y}\|\|_{\psi_{2}}\leq\epsilon_{y}$. Suppose we get $n$ independent observations $x^{(i)},y^{(i)}$ of $x$ and $y$. Assume that $\sigma_{\min}(\sum_{i=1}^{n}x^{(i)}(x^{(i)})^{\top})=\Omega(\beta^{2}n)$. Consider the minimum Frobenius norm solution
@@
Then, there exists an absolute constant $c>0$, such that if $n\geq c(d_{1}+d_{2}+\log(1/p))$, with probability at least $1-p$,
$\|\hat{A}-A^{\ast}\|_{2}=\;$ $\mathcal{O}\big{(}\beta^{-1}\big{(}(\epsilon_{x}+\epsilon_{y})\log^{1/2}(n/p)+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}\big{)}\big{)}.$

### v3

@@ 4.3 Matrix factorization bound
Given two $m\times n$ matrices $A,B$, we are interested in bounding $\min_{S^{\top}S=I}\|SA-B\|_{F}$ using $\|A^{\top}A-B^{\top}B\|_{F}$. The minimum problem is known as the orthogonal Procrustes problem, solved in Schoenemann (1964); Schönemann (1966). Specifically, the minimum is attained at $S=UV^{\top}$, where $U\Sigma V^{\top}=BA^{\top}$ is its singular value decomposition.
If $m\leq n$ and $\mathrm{rank}(A)=m$, then the following lemma from Tu et al. (2016) establishes that the distance between $A$ and $B$ is of the same order of $\|A^{\top}A-B^{\top}B\|_{F}$.
Lemma 3 (Tu et al. (2016, Lemma 5.4)).
@@
$\overset{(i)}{=}\|A\|_{F}^{2}+\|B\|_{F}^{2}-2\mathrm{tr}(\Sigma)$
@@
$\overset{(ii)}{=}\;$ $\sum\nolimits_{i=1}^{d}\sigma_{i}(A^{\top}A+B^{\top}B)-2\sum\nolimits_{i=1}^{d^{\prime}}\sigma_{i}(BA^{\top}),$
@@
4.4 Perturbed linear regression bound
Identifying the dynamics of the latent model requires solving linear regression (3.4). A standard assumption in analyzing linear regression $y=A^{\ast}x+e$ is that $\mathrm{Cov}(x)$ has full rank. However, as discussed in §3.1, we need to handle rank-deficient $\mathrm{Cov}(x)$ in the first $\ell$ steps of system identification. Moreover, the latent state estimates $\hat{z}_{t}$ contain errors. Both issues are addressed in the following lemma.
Lemma 5 (Perturbed rank-deficient linear regression).
Define random vector $y^{\ast}:=A^{\ast}x^{\ast}+e$, where $x^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$ and $e\sim\mathcal{N}(0,\Sigma_{e})$ are $d_{1}$ and $d_{2}$ dimensional Gaussian random vectors, respectively. Define $x:=x^{\ast}+\delta_{x}$ and $y:=y^{\ast}+\delta_{y}$ where the perturbation vectors $\delta_{x}$ and $\delta_{y}$ can be correlated with $x^{\ast}$ and $y^{\ast}$. Assume that $\|A^{\ast}\|_{2}$, $\|\Sigma_{\ast}\|_{2}$ and $\|\Sigma_{e}\|_{2}$ are $\mathcal{O}(1)$. Let $\sigma_{\min}^{+}(\Sigma_{\ast}^{1/2})\geq\beta>0$. Suppose we get $n$ observations $x^{(i)}$ and $y^{(i)}$ of $x$ and $y$, where $((x^{\ast})^{(i)})_{i=1}^{n}$ are independent and $(\delta_{x}^{(i)},\delta_{y}^{(i)})_{i=1}^{n}$ can be correlated. Assume that $\sigma_{\min}^{+}(\sum_{i=1}^{n}x^{(i)}(x^{(i)})^{\top})\geq\theta^{2}n$ for some $\theta>0$ that has at least $n^{-1/4}$ dependence on $n$, and that $\|\sum_{i=1}^{n}\delta_{x}^{(i)}(\delta_{x}^{(i)})^{\top}\|_{2}=\mathcal{O}(\epsilon_{x}^{2}n)$, $\|\sum_{i=1}^{n}\delta_{y}^{(i)}(\delta_{y}^{(i)})^{\top}\|_{2}=\mathcal{O}(\epsilon_{y}^{2}n)$ for $\epsilon_{x},\epsilon_{y}>0$. Consider the minimum Frobenius norm solution
@@ 4.4 Perturbed linear regression bound
Then, there exists an absolute constant $c>0$, such that if $n\geq c(d_{1}+d_{2}+\log(1/p))$, with probability at least 1p1-p,
$\|(\hat{A}-A^{\ast})\Sigma_{\ast}^{1/2}\|_{2}=\mathcal{O}(\beta^{-1}(1+\theta^{-1}(\epsilon_{x}+\epsilon_{y}))\epsilon_{x}+\epsilon_{y}+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}).$
@@
$\hat{A}DG^{\top}GD^{\top}+\hat{A}(DG^{\top}\Delta_{x}+\Delta_{x}^{\top}GD^{\top}+\Delta_{x}^{\top}\Delta_{x})$
$=\;$ $A^{\ast}DG^{\top}GD^{\top}+A^{\ast}DG^{\top}\Delta_{x}+(E^{\top}+\Delta_{y}^{\top})(GD^{\top}+\Delta_{x}).$
@@
$=\;$ $A^{\ast}DG^{\top}\Delta_{x}-\hat{A}(DG^{\top}\Delta_{x}+\Delta_{x}^{\top}GD^{\top}+\Delta_{x}^{\top}\Delta_{x})+(E^{\top}+\Delta_{y}^{\top})(GD^{\top}+\Delta_{x}).$
@@
$\|(\hat{A}-A^{\ast})D\|_{2}=$ $\big{\|}\big{(}A^{\ast}DG^{\top}\Delta_{x}-\hat{A}(DG^{\top}\Delta_{x}+\Delta_{x}^{\top}GD^{\top}+\Delta_{x}^{\top}\Delta_{x})$
$\quad+(E^{\top}+\Delta_{y}^{\top})(GD^{\top}+\Delta_{x})\big{)}(D^{\top})^{\dagger}(G^{\top}G)^{-1}\big{\|}_{2}.$
For $\|\hat{A}\|_{2}$, we make the following claim, whose proof is deferred to §4.4.1.
Claim 1.
As long as $n\geq 16(d_{1}+d_{2}+\log(1/p))$, with probability at least 14p1-4p,
$\|\hat{A}\|_{2}=\mathcal{O}\big{(}\|A^{\ast}\|_{2}(1+\theta^{-1}\epsilon_{x})+\theta^{-1}\epsilon_{y}\big{)}.$
The proof of Claim 1 is deferred to §4.4.1, where we analyze the inversion of $X^{\top}X$. Note that we may also bound $\|\hat{A}\|_{2}$ by analyzing the inversion of $DG^{\top}GD^{\top}$, potentially without the requirement on $\theta$, but this requires stronger condition on $\epsilon_{x}$ and $\epsilon_{y}$, and the bound is not directly applicable to the minimum Frobenius norm solution. The requirement on $\theta$ remains beneficial for numerical stability.
Then, since $\|D^{\dagger}\|_{2}=(\sigma_{\min}^{+}(\Sigma_{\ast}^{1/2}))^{-1}\leq\beta^{-1}$,
$\|(\hat{A}-A^{\ast})D\|_{2}=\;$ $\mathcal{O}(\beta^{-1}\|\hat{A}\|_{2})\big{(}\|D\|_{2}\|G\|_{2}\|\Delta_{x}\|_{2}+\|\Delta_{x}\|_{2}^{2})\|(G^{\top}G)^{-1}\|_{2}\big{)}$
@@
By Wainwright (2019, Theorem 6.1), the Gaussian ensemble $G$ satisfies that with probability at least 1p1-p,
@@
$\geq\;$ $n^{1/2}-d_{1}^{1/2}-(2\log(1/p))^{1/2}).$
Since $n\geq 8d_{1}+16\log(1/p)$, we have $\|G\|_{2}=\mathcal{O}(n^{1/2})$ and $\sigma_{\min}(G)=\Omega(n^{1/2})$. It follows that $\|(G^{\top}G)^{-1}\|_{2}=\mathcal{O}(n^{-1})$ and $\|G^{\dagger}\|_{2}=\mathcal{O}(n^{-1/2})$. Similarly, $\|E\|_{2}=\mathcal{O}((\|\Sigma_{e}\|_{2}n)^{1/2})$. Note that $\Delta_{x}^{\top}\Delta_{x}=\sum_{i=1}^{n}\delta_{x}^{(i)}(\delta_{x}^{(i)})^{\top}$. By our assumption, $\|\Delta_{x}\|_{2}=\mathcal{O}(\epsilon_{x}n^{1/2})$. Similarly, $\|\Delta_{y}\|_{2}=\mathcal{O}(\epsilon_{y}n^{1/2})$. Hence,
$\|(\hat{A}-A^{\ast})D\|_{2}=\;$ $\mathcal{O}(\beta^{-1}\|\hat{A}\|_{2})(\|D\|_{2}\epsilon_{x}+\epsilon_{x}^{2})+\mathcal{O}(1)(\|G^{\dagger}E\|_{2}+\epsilon_{y}+\beta^{-1}(\|\Sigma_{e}^{1/2}\|_{2}+\epsilon_{y})\epsilon_{x})$
$=\;$ $\mathcal{O}(\beta^{-1}\|\hat{A}\|_{2}\epsilon_{x}+\epsilon_{y}+\|G^{\dagger}E\|_{2}),$
where we consider $\epsilon_{x}$ and $\epsilon_{y}$ as quantities much smaller than one such that terms like $\epsilon_{x}^{2},\epsilon_{x}\epsilon_{y}$ are absorbed into $\epsilon_{x},\epsilon_{y}$. It remains to control $\|G^{\dagger}E\|_{2}$. Mhammedi et al. (2020, Section B.2.11) prove via a covering number argument that with probability at least 1p1-p,
@@
$=\;$ $\mathcal{O}(\beta^{-1}\|\hat{A}\|_{2}\epsilon_{x}+\epsilon_{y}+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2})$
$=\;$ $\mathcal{O}(\beta^{-1}(1+\theta^{-1}(\epsilon_{x}+\epsilon_{y}))\epsilon_{x}+\epsilon_{y}+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}),$
@@
4.4.1 Proof of Claim 1
The minimum Frobenius norm solution $\hat{A}$ is given by the following closed-form expression in terms of the pseudoinverse (Moore-Penrose inverse):
$\hat{A}$=\; $Y^{\top}X(X^{\top}X)^{\dagger}$ (4.6)
$=\;$ $(A^{\ast}(X^{\ast})^{\top}+E^{\top}+\Delta_{y}^{\top})X(X^{\top}X)^{\dagger}$
$=\;$ $A^{\ast}(X^{\dagger}X^{\ast})^{\top}+(X^{\dagger}E)^{\top}+(X^{\dagger}\Delta_{y})^{\top}$
$=\;$ $A^{\ast}(X^{\dagger}X)^{\top}-A^{\ast}(X^{\dagger}\Delta_{x})^{\top}+(X^{\dagger}E)^{\top}+(X^{\dagger}\Delta_{y})^{\top}.$
@@ 4.4.1 Proof of Claim 1
By our assumption,
$\|\Delta_{x}\|_{2}=\mathcal{O}(\epsilon_{x}n^{1/2}),\quad\|\Delta_{y}\|_{2}=\mathcal{O}(\epsilon_{y}n^{1/2}).$
Similar to the proof of Lemma 5, by Mhammedi et al. (2020, Section B.2.11), with probability at least 1p1-p,
@@
$\|\hat{A}\|_{2}\leq\;$ $\|A^{\ast}\|_{2}+\mathcal{O}(\theta^{-1}n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2})+\mathcal{O}((\|A^{\ast}\|_{2}\epsilon_{x}+\epsilon_{y})n^{1/2}\theta^{-1}n^{-1/2})$
$=\;$ $\mathcal{O}(\|A^{\ast}\|_{2}(1+\theta^{-1}\epsilon_{x})+\theta^{-1}\epsilon_{y})+\mathcal{O}(\theta^{-1}n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}).$
Hence, as long as $\theta$ has at least $n^{-1/4}$ dependence on $n$,
$\|\hat{A}\|_{2}=\;$ $\mathcal{O}\big{(}\big{(}\|A^{\ast}\|_{2}(1+\theta^{-1}\epsilon_{x})+\theta^{-1}\epsilon_{y}\big{)}\big{)}.$
In Lemma 5, if $\Sigma_{\ast}$ has full rank and $\lambda_{\min}(\Sigma_{\ast})\geq\beta>0$ and $\lambda_{\min}(\sum_{i=1}^{n}x^{(i)}(x^{(i)})^{\top})=\Omega(\beta^{2}n)$, then
@@
Lemma 6 (Perturbed linear regression).
Define random variable $y^{\ast}=A^{\ast}x^{\ast}+e$, where $x^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$ and $e\sim\mathcal{N}(0,\Sigma_{e})$ are $d_{1}$ and $d_{2}$ dimensional random vectors. Assume that $\|A^{\ast}\|_{2}$, $\|\Sigma_{\ast}\|_{2}$ and $\|\Sigma_{e}\|_{2}$ are $\mathcal{O}(1)$, and $\lambda_{\min}(\Sigma_{\ast}^{1/2})\geq\beta>0$. Define $x:=x^{\ast}+\delta_{x}$ and $y:=y^{\ast}+\delta_{y}$ where the perturbation vectors $\delta_{x}$ and $\delta_{y}$ can be correlated with $x^{\ast}$ and $y^{\ast}$. Suppose we get $n$ independent observations $x^{(i)},y^{(i)}$ of $x$ and $y$. Assume that $\lambda_{\min}(\sum_{i=1}^{n}x^{(i)}(x^{(i)})^{\top})=\Omega(\beta^{2}n)$, $\|\sum_{i=1}^{n}\delta_{x}^{(i)}(\delta_{x}^{(i)})^{\top}\|_{2}=\mathcal{O}(\epsilon_{x}^{2}n)$, $\|\sum_{i=1}^{n}\delta_{y}^{(i)}(\delta_{y}^{(i)})^{\top}\|_{2}=\mathcal{O}(\epsilon_{y}^{2}n)$ for $\beta,\epsilon_{x},\epsilon_{y}>0$. Consider the minimum Frobenius norm solution
@@
Then, there exists an absolute constant $c>0$, such that if $n\geq c(d_{1}+d_{2}+\log(1/p))$, with probability at least 1p1-p,
$\|\hat{A}-A^{\ast}\|_{2}=\mathcal{O}\big{(}\beta^{-1}\big{(}\epsilon_{x}+\epsilon_{y}+n^{-1/2}(d_{1}+d_{2}+\log(1/p))^{1/2}\big{)}\big{)}.$